# Renormalon-like factorial enhancements to power expansion/OPE expansion in super-renormalizable QFTs

### v1

@@ Abstract
In this work, we investigate the power-expansion/OPE expansion in super-renormalizable QFT. In comparison to “marginal” short distance asymptotics, the power-expansion in super-renormalizable theory has a higher-level of correlation with the expansion in the coupling constants, and the fixed-point perturbation theory for the coefficient functions suffers increasing level of IR divergences when the order of power-expansion/coupling-constant expansion increases. This increasing IR sensitivity anticipates the presence of vacuum condensates of non-trivial local operators in the OPE expansion, which not only cancel the IR divergences for coefficient functions through their UV divergences, but also shape the structure of logarithms at each power in the power-expansion. This UV-IR relationship has been used to argue for the renormalon cancellation between coefficient functions and operators in marginal asymptotics, where the residue IR sensitivity of coefficient functions to be canceled by the operators manifests through high-order factorial growths after renormalization subtraction.
However, in supper-renomalizable theory, high-order behaviors in coupling constants in the coefficient functions are simultaneously high-power behaviors in the power expansion, and one may wounder if there are also factorial growths for high-order/high-power terms in the coefficient functions, as they also require large-numbers of IR subtractions. In this work, we address the issue regarding the high-power behavior of power-expansion/OPE-expansion in supper-renormalizable theory. Using an $O(N)$-model with $N$-components scalars coupled through quartic interaction at the next-to-leading $\frac{1}{N}$ order in the large-$N$ expansion, we show that the IR subtractions indeed cause addition factorial enhancements for high-power terms in the coefficient functions. Moreover, there are also factorial enhancements for the operator condensates, and the factorial enhancements cancel between coefficient functions and operators only off-diagonally across different powers. The factorial enhancements can be both alternating and non-alternating. The former are similar to “UV renormalon” of coefficient functions and cancel with factorial enhancements of operators at lower powers in diagrams with negative degrees of UV divergences. The later are similar to “IR renormalon” and cancel with factorial enhancements of renormalized operators at higher powers in diagrams with positive degrees of UV divergences. The factorial enhancement itself will render the momentum-space power expansion divergent.
@@ 1 Introduction
Operator product expansion (OPE) Wilson (1971) is among one of the major properties concerning short distance behavior of correlation functions in local QFTs. In the “simplest” local QFTs, the conformal field theories (CFTs), existence of OPE in the sense of asymptotic expansion is sufficient to imply the absolute convergence of OPE in the coordinate-space analyticity region Luscher and Mack (1975); Mack (1977); Kravchuk et al. (2021). In massive QFTs with CFTs as their short distance fixed-points, on the other hand, although still exist at worst in the sense of asymptotic expansion in the Euclidean region, the behavior of OPE is much less controlled due to the existence of non-trivial IR physics. Not only derivative operators are much harder to be resumed, the coefficient functions for individual operators are no-longer homogeneous functions and will contain non-trivial $z^{2}m^{2}$ dependencies Novikov et al. (1985); Lukyanov and Zamolodchikov (1997) that couple the short distance and large distance scales.
@@
In super-renormalizable QFTs, namely, QFTs that are obtained by perturbing the short-distance fixed points through relevant operators, the $z^{2}m^{2}$ dependencies in the coefficient functions can still be separated into well-defined powers and calculated order by order in fixed-point perturbation theory (also called CFT perturbation theory) with appropriate IR subtractions Chetyrkin et al. (1982); Chetyrkin (1983). The IR subtractions introduce scale-scheme dependencies to the coefficient functions that are canceled power-wisely in the OPE expansion, by the scale-scheme dependencies introduced to the operators during their UV-divergence renormalization. To some sense, the IR sensitivity in the fixed-point perturbation theory anticipates the existence of non-trivial operator condensates Novikov et al. (1980); Shifman et al. (1979); David (1984, 1986); Novikov et al. (1985); Mueller (1985); Ji (1995); Beneke (1999); Beneke et al. (1998); Braun et al. (2004); Shifman (2015); Dunne et al. (2015). However, exactly due to the large-numbers of IR subtractions required to define the high-power terms in the coefficient functions, their quantitative behaviors can be significantly modified in comparison to quantities defined through absolutely-convergent Feynman integrals. In the worst scenario, factorial growths can be developed at a relatively small numbers of diagrams and modify the convergence property of the OPE expansion or at least, the convergence of power-expansion for the coefficient functions.
@@
It is the task of this work to investigate the issue of high-power behavior of coefficient functions in super-renormalizable QFTs in the presence of large numbers of IR subtractions. Using the large-$N$ expansion of an $O(N)$ model with quartic interaction, one can remove the factorial growths in the numbers of Feynman diagrams and focus on the renormalization effects. We show that the IR-subtraction-introduced factorial growths in the power expansion in the Euclidean momentum space indeed exist and are canceled between coefficient functions and operators off-diagonally across different powers. The off-diagonal nature of the cancellation implies that the power expansion in momentum space is divergent. To make the OPE expansion convergent, one must subtract all the factorial growths from operators and add back to coefficient functions. This requires infinitely-many redefinitions of operators and can not be achieved through a single “Borel prescription”. On the other-hand, in the coordinate space, due the overall $\frac{1}{(n!)^{2}}$ suppression, the $n!$ introduced by the IR subtraction is insufficient to modify the convergence properties of the power-expansion/OPE expansion.
@@
    In the Sec. 2, we provide a worked example of the OPE calculation up to all the non-trivial coefficient functions up to ${\cal O}(z^{4})$, for the scalar two-point function in the 2D one-component $\phi^{4}$ theory. The section is largely for pedagogical purpose to convince unfamiliar readers the fact that the coefficient functions in super-renormalized QFTs can still be calculated through IR-renormalized massless Feynman integrals regulated in dimensional regularization (DR). It will also demonstrates how the UV-IR relationship between coefficient functions and operators works in this example.
@@
    In the Sec. 3, we start to investigate the high-power behavior of IR renormalized coefficient functions in the large-$N$ model. We first perform the calculation through IR renormalized massless bubble-chain diagrams, for the diagrams with zero and one mass insertions (but with $n$ and $n-1$ massless bubbles). The $n!$ enhancement is generated through a novel $\frac{1}{\epsilon}\times\epsilon$ effect that is absent in 4D bubble-chain calculations. We also show how the $n!$ enhancements cancel with that of operators in a way without directly calculating the massive bubble-chains. We then double-check with the short-distance calculation by introducing Mellin-Barnes representations for the full massive bubble chain diagram and performing the power-expansion by shifting the contours for the Mellin-Barnes. The result completely agrees with the short-distance calculation and allows one to effectively obtain the leading large-$n$ asymptotics for the coefficient function at the power $\frac{1}{(p^{2})^{n+2}}$ after summing over all possible numbers of mass insertions at the given power.
@@ 2 OPE in supper-renormalizable theory: $\phi^{4}_{2}$ as an example
In this section we provide a worked example for OPE calculation both in coordinate space and momentum space, in the massive $\phi^{4}_{2}$ theory for the scalar two-point function. We always stay in Euclidean signature in this work. Familiar readers can skip this section.
@@
Throughout the paper we use DR with $D=2-2\epsilon$ for all the UV/IR divergences and $\mu_{0}^{2}=\mu^{2}\frac{e^{\gamma_{E}}}{4\pi}$ are the standard renormalization scales. Notice that the Wick-ordering is with respect to the free Gaussian field with mass $m$. This defines the mass parameter $m$, as well as the bare mass
@@
In the $\epsilon\rightarrow 0$ limit, all correlation functions of the $\phi$ field at separated points are finite and can be calculated directly at $D=2$ by neglecting the one-loop tadpoles.
@@
in this theory, up to all the coefficient functions contributing to ${\cal O}(z^{4})$. Since in the coefficient functions there are no inverse power in $g$, only $g^{2}$ (two-loop) is required for this calculation. To show that the OPE is correct, it is helpful to show that the coefficient functions combined with the operator condensates reproduces the correct small-$z^{2}$ expansion of tree-level and two-loop massive diagrams of $S$. In the appendix, we show using explicit calculation that for the tree-level and two-loop two-point function
@@
Figure 3: The coefficient functions for various operators at the order ${\cal O}(z^{0})$ and ${\cal O}(z^{2})$. The dashed lines are massless propagators. The big black circles are insertion of the bare mass $-m_{0}^{2}$. The double lines denote the “external soft fields” from condensates attached to the coefficient functions. The tree-level coefficient $1$ for $\phi^{2}$ and $-\frac{1}{2d}$ for $\phi(i\partial)^{2}\phi$ are omitted in the figure, but are crucial for coordinates space expansion.
@@ 2.1 Leading power ($z^{0}$) and next-to-leading power ($z^{2}$)
At leading power, in momentum space there is only a single term $\frac{1}{p^{2}}$ [^1] Notice that the momentum-space two point function in $D$-dimension in our convention has an overall factor $\mu_{0}^{2-D}$. This factor will not participate in the renormalization and will be omitted in this work.. In coordinate space, on the other hand, there is already a log
@@
$C_{0}^{(0)}(z^{2}\mu^{2},d)+\langle\phi^{2}\rangle_{0}\bigg{|}_{d\rightarrow 2}\rightarrow-\frac{1}{4\pi}\ln\frac{z^{2}m^{2}e^{2\gamma_{E}}}{4}\ .$ (18)
@@
Clearly, the $\frac{1}{4\pi\epsilon}$ of the operator renormlaization exactly cancels the IR divergence of the bare coefficient function. Moreover, the above allows the generalization to include high-order corrections to the operator condensates:
@@
Thus, the $\mu^{2}$ determines the separation point between short distance and large distance fluctuations in the $\overline{\rm MS}$ scheme. The $\langle:\phi^{2}:\rangle$ start to contribute at $g^{2}$ and is given in Eq. (12). It exactly reproduces the $\zeta_{3}$ term at $z^{0}$ in the first line of Eq. (2). Higher order corrections will not contribute any new logarithms at leading power, but only modifies this condensate. Notice that the bare coefficient function completely determines the $\frac{1}{4\pi\epsilon}$ for the operator renormalization.
@@
There are two condensates $\phi^{2}$ and $\phi(i\partial)^{2}\phi$ in the OPE. The bare coefficient functions follows from figure 3
@@
$S^{(2)}(p^{2})(p^{2})^{2}=-m_{0}^{2}-\frac{g}{2}\langle\phi^{2}\rangle_{0}\ .$ (27)
This determines the renormalization of the bare mass
@@
Here we have introduced short distance renormalized mass $m_{r}^{2}$ that is $\epsilon$ dependent, but is finite in the $\epsilon\rightarrow 0$ limit. It can be regarded as the bare mass subtracted only the short-distance contribution to the mass rernomalization, but without including any effects at the IR scale $m$. Due to this, it is the most natural mass that appears in the short distance calculation and in the renormalized coefficient functions. Here, we have seen that by requiring the IR finiteness of renormalized coefficient functions automatically determines the mass renormlaization.
To determine the renormalization of $\phi\partial^{2}\phi$, one can write
$\langle\phi(i\partial)^{2}\phi\rangle_{r}=a_{1}m_{r}^{2}+a_{2}g\langle\phi^{2}\rangle_{r}+\langle\phi(i\partial)^{2}\phi\rangle_{r}\ ,$ (29)
with $a_{1},a_{2}$ are $\frac{1}{\epsilon}$ poles. Now, requiring finiteness of renormalized coefficient functions
$C_{{\rm I},r}^{(2)}=C_{0}^{(2)}+\frac{1}{4\pi\epsilon}C_{\phi^{2}}^{(2)}-\frac{z^{2}}{2d}m_{r}^{2}a_{1}\rightarrow{\cal O}(1)\ ,$ (30)
@@
This uniquely determines the operator renormalization
@@
The above is consistent with $m_{0}^{2}=m^{2}-\frac{g}{2}I_{0}$ and $I_{0}=I_{r}+\frac{1}{4\pi\epsilon}$. Thus, the renormalization constants can also be read from the above equation and agrees with the one determined from the IR-UV cancellation requirement. It is easy to show that the above coefficient functions combined with the tree-level condensates exactly reproduces the full expansion at $z^{2}$ and ${\cal O}(g^{0})$. Furthermore, from the above the RGE for the renormalized operator $\phi\partial^{2}\phi$ and the general solution is

### v2

@@ Abstract
In this work, we investigate the power-expansion/OPE expansion in super-renormalizable QFTs. In comparison to “marginal” short distance asymptotics, the power-expansion in super-renormalizable theory has a higher level of correlation with the coupling constant expansion, and the fixed-point perturbation theory for the coefficient functions suffers increasing level of IR divergences when the order of power-expansion/coupling constant expansion increases. This increasing IR sensitivity anticipates the presence of vacuum condensates of non-trivial local operators in the OPE expansion, which not only cancel the IR divergences for coefficient functions through their UV divergences, but also shape the structure of logarithms at each power in the power-expansion. This UV-IR correspondence has been used to argue for the renormalon cancellation between coefficient functions and operators in marginal asymptotics, where the residue IR sensitivity of coefficient functions to be canceled by the operators manifests through high-order factorial growths after renormalization subtraction.
However, in supper-renomalizable theory, high-order behavior in coupling constants in the coefficient functions is simultaneously large-order behavior in the power expansion, and one may wonder if there are also factorial growths for high-order/power terms in the coefficient functions, as they also require large-numbers of IR subtractions. In this work, we address the issue regarding the asymptotic behavior of power-expansion/OPE-expansion in supper-renormalizable theory. Using an $O(N)$-model with $N$-components scalars coupled through quartic interaction at the next-to-leading $\frac{1}{N}$ order in the large-$N$ expansion, we show that the IR subtractions indeed cause addition factorial enhancements for large order/power terms in the coefficient functions. Moreover, there are also factorial enhancements for the operator condensates, and the factorial enhancements cancel between coefficient functions and operators only off-diagonally across different powers. The factorial enhancements can be both alternating and non-alternating. The former is similar to the “UV renormalon” of coefficient functions and cancels with factorial enhancements of operators at lower powers in diagrams with negative degrees of UV divergences. The latter is similar to “IR renormalon” and cancels with factorial enhancements of renormalized operators at higher powers in diagrams with positive degrees of UV divergences. The factorial enhancement itself will render the momentum-space power expansion divergent.
@@ 1 Introduction
Operator product expansion (OPE) Wilson (1971) is among one of the major properties concerning short distance behavior of correlation functions in local QFTs. In the “simplest” local QFTs, the conformal field theories (CFTs), the existence of OPE in the sense of asymptotic expansion is sufficient to imply the absolute convergence of OPE in the coordinate-space analyticity region Luscher and Mack (1975); Mack (1977); Kravchuk et al. (2021). In massive QFTs with CFTs as their short-distance fixed-points, on the other hand, although still exist at worst in the sense of asymptotic expansion in the Euclidean region, the behavior of OPE is much less controlled due to the existence of non-trivial IR physics. Not only derivative operators are much harder to be resumed, the coefficient functions for individual operators are no longer homogeneous functions and will contain non-trivial $z^{2}m^{2}$ dependencies Novikov et al. (1985); Lukyanov and Zamolodchikov (1997) that couple the short distance and large distance scales.
@@
In super-renormalizable QFTs, namely, QFTs that are obtained by perturbing the short-distance fixed points through relevant operators, the $z^{2}m^{2}$ dependencies in the coefficient functions can still be separated into well-defined powers and calculated order by order in fixed-point perturbation theory (also called CFT perturbation theory) with appropriate IR subtractions Chetyrkin et al. (1982); Chetyrkin (1983). The IR subtractions introduce scale-scheme dependencies to the coefficient functions that are canceled power-wisely in the OPE expansion, by the scale-scheme dependencies introduced to the operators during their UV-divergence renormalization. To some sense, the IR sensitivity in the fixed-point perturbation theory anticipates the existence of non-trivial operator condensates Novikov et al. (1980); Shifman et al. (1979); David (1984, 1986); Novikov et al. (1985); Mueller (1985); Ji (1995); Beneke (1999); Beneke et al. (1998); Braun et al. (2004); Shifman (2015); Dunne et al. (2015). However, exactly due to the large numbers of IR subtractions required to define the high-order terms in the coefficient functions, their quantitative behaviors can be significantly modified in comparison to quantities defined through absolutely convergent Feynman integrals. In the worst scenario, factorial growths can be developed at a relatively small number of diagrams and modify the convergence property of the OPE expansion or at least, the convergence of power-expansion for the coefficient functions.
@@
It is the task of this work to investigate the issue of asymptotic behavior of coefficient functions in super-renormalizable QFTs in the presence of large numbers of IR subtractions. Using the large-$N$ expansion of an $O(N)$ model with quartic interaction, one can remove the factorial growths in the numbers of Feynman diagrams and focus on the renormalization effects. We show that the IR-subtraction-introduced factorial growths in the power expansion in the Euclidean momentum space indeed exist and are canceled between coefficient functions and operators off-diagonally across different powers. The off-diagonal nature of the cancellation implies that the power expansion in momentum space is divergent. To make the OPE expansion convergent, one must subtract all the factorial growths from operators and add back to coefficient functions. This requires infinitely many redefinitions of operators and can not be achieved through a single “Borel prescription”. On the other hand, in the coordinate space, due to the overall $\frac{1}{(n!)^{2}}$ suppression, the $n!$ introduced by the IR subtraction is insufficient to modify the convergence properties of the power-expansion/OPE expansion.
@@
    In the Sec. 2, we provide a worked example of the OPE calculation up to all the non-trivial coefficient functions up to ${\cal O}(z^{4})$, for the scalar two-point function in the 2D one-component $\phi^{4}$ theory. The section is largely for pedagogical purposes to convince unfamiliar readers the fact that the coefficient functions in super-renormalized QFTs can still be calculated through IR-renormalized massless Feynman integrals regulated in dimensional regularization (DR). It will also demonstrate how the UV-IR correspondence between coefficient functions and operators works in this example.
@@
    In the Sec. 3, we start to investigate the asymptotic behavior of IR renormalized coefficient functions in the large-$N$ model. We first perform the calculation through IR renormalized massless bubble-chain diagrams, for the diagrams with zero and one mass insertions (but with $n$ and $n-1$ massless bubbles). The $n!$ enhancement is generated through a novel $\frac{1}{\epsilon}\times\epsilon$ effect that is absent in 4D bubble-chain calculations. We also show how the $n!$ enhancements cancel with that of operators in a way without directly calculating the massive bubble-chains. We then double-check with the short-distance calculation by introducing Mellin-Barnes representations for the full massive bubble chain diagram and performing the power-expansion by shifting the contours for the Mellin-Barnes. The result completely agrees with the short-distance calculation and allows one to effectively obtain the leading large-$n$ asymptotics for the coefficient function at the power $\frac{1}{(p^{2})^{n+2}}$ after summing over all possible numbers of mass insertions at the given power. Our work therefore provides a concrete example showing the “OPE divergence” phenomenon Shifman (1994) as asked in Marino et al. (2023).
@@ 2 OPE in supper-renormalizable theory: $\phi^{4}_{2}$ as an example
In this section, we provide a worked example for OPE calculation both in coordinate and momentum spaces, in the massive $\phi^{4}_{2}$ theory for the scalar two-point function. We always stay in Euclidean signature in this work. Although this section may look familiar to experts, the author still suggests reading this section, which could also serve as a motivation for later sections.
@@
Throughout the paper we use DR with $D=d=2-2\epsilon$ for all the UV/IR divergences and $\mu_{0}^{2}=\mu^{2}\frac{e^{\gamma_{E}}}{4\pi}$ are the standard renormalization scales. Notice that the Wick-ordering is with respect to the free Gaussian field with mass $m$. This defines the mass parameter $m$, as well as the bare mass
@@
In the $\epsilon\rightarrow 0$ limit, all correlation functions of the $\phi$ field at separated points are finite and can be calculated directly at $D=2$ by neglecting the one-loop tadpoles. Notice that the parameters $m$ and $g$ are $\epsilon$ independent.
@@
in this theory, up to all the coefficient functions contributing to ${\cal O}(z^{4})$. Since in the coefficient functions there are no inverse powers in $m^{2}$ and $g$, only $g^{2}$ (two-loop) is required for this calculation. To show that the OPE is correct, it is helpful to show that the coefficient functions combined with the operator condensates reproduce the correct small-$z^{2}$ expansion of tree-level and two-loop massive diagrams of $S$. In the appendix A, we show using explicit calculations that for the tree-level and two-loop two-point function
@@
Figure 3: The coefficient functions for various operators at the order ${\cal O}(z^{0})$ and ${\cal O}(z^{2})$. The dashed lines are massless propagators. The big black circles are insertion of the bare mass $-m_{0}^{2}$. The double lines denote the “external soft fields” from condensates attached to the coefficient functions. They will be amputated in the coefficient function. The tree-level coefficient $1$ for $\phi^{2}$ and $-\frac{1}{2d}$ for $\phi(i\partial)^{2}\phi$ are omitted in the figure, but are crucial for coordinates space expansion.
@@ 2.1 Leading power ($z^{0}$) and next-to-leading power ($z^{2}$)
At leading power, in the momentum space there is only a single term $\frac{1}{p^{2}}$ [^1] Notice that the momentum-space two point function in $D$-dimension in our convention has an overall factor $\mu_{0}^{2-D}$. This overall factor will not participate in the momentum-space renormalization and will be omitted in this work. In the coordinate space it participates in the momentum integral and can not be neglected.. In coordinate space, on the other hand, there is already a log
@@
$C_{\rm I}^{(0)}(z^{2}\mu^{2},d)+\langle\phi^{2}\rangle_{0}\bigg{|}_{d\rightarrow 2}\rightarrow-\frac{1}{4\pi}\ln\frac{z^{2}m^{2}e^{2\gamma_{E}}}{4}\ .$ (18)
@@
Clearly, the $\frac{1}{4\pi\epsilon}$ pole of the bare operator exactly cancels the IR divergence of the bare coefficient function, and it can also be regarded as the IR subtraction or IR counter-term for the coefficient function. Moreover, the above allows the generalization to include high-order corrections to the operator condensates:
@@
Thus, the $\mu^{2}$ determines the separation point between short distance and large distance fluctuations in the $\overline{\rm MS}$ scheme. The $\langle:\phi^{2}:\rangle$ start to contribute at $g^{2}$ and is given in Eq. (12). It exactly reproduces the $\zeta_{3}$ term at $z^{0}$ in the first line of Eq. (2). Higher order corrections will not contribute any new logarithms at the leading power, but only modifies this condensate. Notice that the bare coefficient function completely determines the $\frac{1}{4\pi\epsilon}$ for the operator renormalization.
@@
There are two condensates $\phi^{2}$ and $\phi(i\partial)^{2}\phi$ in the OPE. The bare coefficient functions follow from figure 3
@@
$S^{(2)}(p^{2})(p^{2})^{2}=-m_{0}^{2}-\frac{g}{2}\langle\phi^{2}\rangle_{0}=-m_{0}^{2}-\frac{g}{8\pi\epsilon}-\frac{g}{2}\langle\phi^{2}\rangle_{r}\ ,$ (27)
where we have used the renormalization constant Eq. (19). By requiring its finiteness, one determines the renormalization of the bare mass
@@
Here we have introduced short distance renormalized mass $m_{r}^{2}$ that is $\epsilon$ dependent, but is finite in the $\epsilon\rightarrow 0$ limit. It can be regarded as the bare mass subtracted only the short-distance contribution to the mass rernomalization, but without including any effects at the IR scale $m$. Due to this, it is the most natural mass that appears in the short distance calculation and in the renormalized coefficient functions.
Here, we have seen that by requiring the IR finiteness of renormalized coefficient functions automatically determines the mass renormlaization. This also resolves the “puzzle” concerning the consistency between the DR rule that throws away scaleless integrals and the fact that the $\overline{\rm MS}$ renormalization is short distance in nature: the mass renormalization in $\phi_{2}^{4}$ comes purely from the one-loop tadpole which becomes scaleless in the massless theory. How this renormalization can be determined without computing any massive integrals? Clearly, the approach here is the way: by requiring IR finiteness of the leading-power coefficient function $C_{\rm I,r}^{(0)}$ uniquely determines the operator renormalization for $\langle\phi^{2}\rangle_{0}$, the latter then determines the mass renormalization by requiring IR finiteness of the next-to-leading power coefficient function. In fact, this pattern will persist to all powers: requiring IR finiteness of the coefficient functions automatically determines the UV poles of operators.
Here we move back to the coordinate space. To determine the renormalization of $\phi\partial^{2}\phi$, one can write
$\langle\phi(i\partial)^{2}\phi\rangle_{0}=a_{1}m_{r}^{2}+a_{2}g\langle\phi^{2}\rangle_{r}+\langle\phi(i\partial)^{2}\phi\rangle_{r}\ ,$ (29)
with $a_{1},a_{2}$ are $\frac{1}{\epsilon}$ poles. Notice that in this paper, we call purely $\frac{1}{\epsilon}$ coefficients such as $a_{1}$, $a_{2}$ that express bare operators in terms of renormalized ones as UV poles of the operators. They are simultaneously IR counter-terms for the coefficient functions. Now, requiring finiteness of renormalized coefficient functions
$C_{{\rm I},r}^{(2)}=C_{\rm I}^{(2)}+\frac{1}{4\pi\epsilon}C_{\phi^{2}}^{(2)}-\frac{z^{2}}{2d}m_{r}^{2}a_{1}\rightarrow{\cal O}(1)\ ,$ (30)
@@
This uniquely determines the operator poles
@@
The above is consistent with $m_{0}^{2}=m^{2}-\frac{g}{2}I_{0}$ and $I_{0}=I_{r}+\frac{1}{4\pi\epsilon}$. Thus, the renormalization constants can also be read from the above equation and agrees with the one determined from the IR-UV cancellation requirement. It is easy to show that the above coefficient functions combined with the tree-level condensates exactly reproduces the full expansion at $z^{2}$ and ${\cal O}(g^{0})$. Furthermore, from the above the RGE for the renormalized operator condensate $\langle\phi\partial^{2}\phi\rangle_{r}$ and the general solution is